# Optimized Multi-Token Joint Decoding with Auxiliary Model for LLM Inference

### v2

@@ 4 Energy Efficiency Analysis
Previous studies (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023; Sun et al., 2023) only focus on the speed of speculative decoding. However, an equally important consideration is energy consumption. To our knowledge, there is no existing work evaluating the impact of speculative decoding on inference energy consumption. Although MTAD and speculative decoding increase the number of FLOPs due to the involvement of a small auxiliary model and the rollback operation, they concurrently reduce the inference time and memory operations, which are key factors of GPU energy consumption (Allen & Ge, 2016; Chen et al., 2011). Consequently, it poses an open question regarding whether speculative decoding increases or decreases overall energy consumption.
@@
where $PW_{flop},PW_{mem}$ denote the power (energy/second) of FLOPs and memory operations, and $T_{flop},T_{mem}$ denote the time spent on these operations. When input batch size increases, $PW_{flop}$ increases until it reaches the power of maximum FLOPs, denoted as $PW^{*}_{flop}$. Meanwhile, $PW_{mem}$ is irrelevant to the input batch size, as it only depends on the memory hardware.
@@
To determine the relative magnitude relationship between $PW_{flop}$ and $PW_{mem}$, we first point out the fact that GPU memory operations in LLM inference are dominated by accessing off-chip global memory, which consumes about 100×100$\times$100 × of energy compared to accessing on-chip shared memory (Jouppi et al., 2021). It is because each multiprocessor on GPU usually has 64KB of on-chip memory shared by multiple threads, while storing a single layer of LLM, say T5-11b (Raffel et al., 2020), requires about 1GB memory. Moreover, Allen and Ge showed that doing sequential read from off-chip memory consumes $20$-$30\%$ more power than running maximum FLOPs (Allen & Ge, 2016). So we have $PW_{mem}>PW^{*}_{flop}\geq PW_{flop}$. Notice that $PW^{*}_{flop}=PW_{flop}$ only if the batch size reaches the maximum parallelization capacity of GPUs. During multinomial sampling and speculative decoding, the batch size is usually small (Leviathan et al., 2023). So most of the computing power is not utilized (Leviathan et al., 2023), which means $PW_{mem}\gg PW_{flop}$.
In addition, previous studies have shown that during LLM inference $T_{mem}\gg T_{flop}$ (Leviathan et al., 2023). Therefore, the energy induced by memory operations, i.e., $PW_{mem}T_{mem}$ dominates $E_{total}$. Since speculative decoding reduces $T_{mem}$ by reducing the number of runs of the large model, it should reduce the inference energy consumption to a similar extent as it reduces time consumption.
To validate our hypothesis, we conducted an experiment to evaluate how batch size influences energy consumption during inference. We ran OPT-13b models on a Nvidia L40 GPUs with 48GB memory. Fixing the total number of runs of the large model while varying the input batch size $b\in\{1,2,4,8,16\}$ for each run, we measured time and energy cost. The details of energy measurement are illustrated in the Appendix D. Table 2 shows the results. As batch size doubles, although the number of FLOPs doubles, the energy consumption per run increases slightly. This observation demonstrates that $PW_{mem}T_{mem}$ dominates $E_{total}$. Moreover, we measured the speed and energy consumption of running multinomial sampling with the large model and speculative decoding using OPT (125M, 13B) and Llama-2 (68M, 13B) models. The results, shown in Table 3, indicate that speculative decoding reduces the energy consumption and the time cost. This observation corroborates our claim to the energy efficiency of speculative decoding.
@@ 5 Experiments
Datasets and Models. We use five public datasets for evaluation: (1) ChatGPT-Prompt (Rashad, 2023), (2) ChatAlpaca (Bian et al., 2023), (3) CNN Dailymail (See et al., 2017), (4) Spider (Yu et al., 2018), and (5) MT-bench (Zheng et al., 2023). Table 8 in the Appendix shows more details of the datasets. Following previous studies (Kim et al., 2023), we use two public LLM families in our experiments: OPT (Zhang et al., 2022) and Llama-2 (Touvron et al., 2023). In this section, we set the large model to be OPT-13B and Llama-2-13B as they are the largest models that can run on a single 40GB GPU, and use Llama-68M (Miao et al., 2023) and OPT-125M as the small models. Appendix C reports additional experiment results with OPT-30B and Llama-2-chat-70B.
Baselines. For each pair of small and large models, we compare our method with four speculative decoding methods: vanilla speculative decoding (speculative(Lee et al., 2018; Chen et al., 2023), Spectr (Sun et al., 2023), SpecInfer (Miao et al., 2023), and BiLD (Kim et al., 2023). Our implementation of MTAD and all the baselines are based on a public implementation of speculative decoding (Bear, 2024). For each method, we let it generate at most 128 tokens for each input and run it for $1,000$ seconds. We open-sourced our code for reproduction. All the methods are stochastic with top-$k$ and top-$p$ sampling. The details of the hyper-parameters (e.g., $k$ and $p$) and machine configurations of the experiments can be found in the Appendix D, E, and F.
@@
[…]
5.1 Comparison with Baselines
Table 4 shows the primary results of our experiments while Table 5 details the block efficiency, defined as the average number of tokens generated per iteration, of different methods. The standard deviations in the tables are computed by repeating each experiment four times [^4] For MT-Bench, the standard deviation also accounts for variations across different tasks.. First, we observe that MTAD is significantly more efficient than all baselines in terms of both energy and time. On one hand, the energy consumption of MTAD is on average 1.54×1.54$\times$1.54 × smaller than that of vanilla speculative decoding. On the other hand, MTAD is 1.101.71×1.10-1.71$\times$1.10 - 1.71 × faster than vanilla speculative decoding, 1.38×1.38$\times$1.38 × faster than BiLD, 1.59×1.59$\times$1.59 × faster than Spectr, and 1.60×1.60$\times$1.60 × faster than SpecInfer. SpecInfer and Spectr has better block efficiency than vanilla speculative decoding, but are slower. This may be due to the fact that they have to verify multiple draft sequences, which introduces extra overhead and may not be perfectly parallelized, especially when the memory overhead exceeds the GPU memory capacity. Meanwhile, MTAD has the best block efficiency without causing any extra overhead, hence it is significantly more efficient.
Llama-2 OPT
spec 2.02$\pm$0.05 2.60$\pm$0.06
BiLD 1.83$\pm$0.10 2.68$\pm$0.36
Spectr 2.73$\pm$0.43 3.45$\pm$0.42
SpecInfer 2.74$\pm$0.46 3.45$\pm$0.40
MTAD 3.17$\pm$0.43 4.30$\pm$0.03
Next, we compare the output perplexity of different algorithms. The perplexity scores of vanilla speculative decoding, SpecInfer, and Spectr are close since their sampling distributions are equivalent. Meanwhile, BiLD approximates the sampling distribution of single-token multinomial sampling but yields better perplexity. It is because we set a strict acceptance threshold for BiLD, which lowers the acceptance rate but ensures every token has a high probability in the large model, thus improving the overall likelihood. More importantly, there is a significant gap between MTAD and other baselines. On average, the perplexity of MTAD is $21.2\%$ lower than that of speculative decoding.
spec MTAD
CD Rouge-L 0.114 0.118
SP EA 11.5 13.0
MT Humanities 2.95 3.15
Extraction 1.80 2.50
Roleplay 3.10 3.80
Math 1.10 1.00
Coding 1.25 1.10
Reasoning 3.80 3.15
STEM 2.85 3.10
Writing 3.80 3.65
Average 2.58 2.68
In addition, to show MTAD indeed improves the downstream effectiveness, we compare the performance metrics of speculative decoding and MTAD on CNNDM, Spider, and MT-Bench datasets. [^5] All speculative baselines are equivalent to multinomial sampling We exclude the other two datasets as they lack explicit downstream metrics. And we exclude the results of OPT models due to their consistently poor performance across all evaluated datasets. As illustrated in Table 6, MTAD outperforms speculative decoding across all three datasets, thereby validating our claim that MTAD achieves superior effectiveness compared to conventional decoding methods that rely on single-token distributions.
5.2 Ablation Study
5.2.1 Number of Beams
First, we investigate how the number of beams used in the beam decoding of the small model affects the inference performance. Table 7 shows the results. Increasing the number of beams improves the quality of the draft tokens, which not only improves the output perplexity but also increases the average acceptance length and hence leads to better efficiency. But we can see that the increment slows down when the number of beams is large enough. In addition, when the number of beams is too large, the inference cost of the small model will become too high.
Table 7: Effect of number of beams to the inference performance on ChatGPT-Prompts dataset.
# beams 2 4 6 8
speed (token/s) $\uparrow$ 55.9 59.9 60.2 61.3
Llama-2 energy (J/token) $\downarrow$ 2.43 2.25 2.22 2.20
perplexity $\downarrow$ 2.44 2.12 2.14 2.10
speed (token/s) $\uparrow$ 51.0 54.1 54.3 55.9
OPT energy (J/token) $\downarrow$ 2.50 2.32 2.36 2.30
perplexity $\downarrow$ 3.63 3.16 3.42 3.19
5.2.2 Acceptance Threshold
Next, we evaluate the effect of acceptance threshold $\tau$. Intuitively, when we increase $\tau$ from 0 to 1, the acceptance criterion becomes more strict, the efficiency drops while the output perplexity increases. Surprisingly, this expectation is only partially correct. As shown in Figure 3, the efficiency indeed drops when $\tau$ increases. However, the perplexity increases when $\tau$ is close to 1. When $\tau=1$, all the draft tokens are rejected, which makes MTAD equivalent to multinomial sampling. Similarly, when $\tau$ is close to 1, the advantage of multi-token joint decoding on effectiveness disappears, hence the perplexity becomes close to the perplexity of multinomial sampling. Another surprising observation is that the perplexity of MTAD is good when $\tau=0$. When $\tau=0$, MTAD is equivalent to generating $\gamma$ tokens using beam decoding with the small model, then generating an additional token with the large model. The fact that MTAD achieves good perplexity when $\tau=0$ can be explained by the fact that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s predictions deviate from those of the large model” (Kim et al., 2023; Leviathan et al., 2023). Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MTAD is relatively stable, suggesting that MTAD is not sensitive to the acceptance threshold.
@@ 6 Related Work
A direction closer to our work is non-autoregressive decoding. It is first proposed by (Gu et al., 2017) to generate multiple tokens in parallel. That is, the model simultaneously predicts $p(x_{t+k}|x_{1:t})$ ($k=1,2,\ldots$). Subsequent studies further improved the performance of parallel decoding by incorporating additional information (Wang et al., 2019; Sun et al., 2019; Li et al., 2019) or using additional iterations to refine predictions (Ghazvininejad et al., 2019; Lee et al., 2018; Guo et al., 2020). However, these works require continuous training of the model and usually either compromise the model effectiveness or require task-dependent techniques to achieve comparable performance (Kim et al., 2023).
Speculative Decoding. Speculative decoding was recently proposed in (Leviathan et al., 2023; Chen et al., 2023) as a way to accelerate LLM inference. Spectr (Sun et al., 2023) enhances speculative decoding by letting the small model generate multiple i.i.d. draft sequences. While speculative decoding and Spectr use the large model to verify all the tokens drafted by the small model, BiLD (Kim et al., 2023) only calls the large model when the probability output by the small model is below a pre-defined threshold $\tau_{1}$. The large model rejects a token if its negative log-likelihood is larger than threshold $\tau_{2}$. SpecInfer (Miao et al., 2023) uses one or multiple small models to generate a draft token tree to increase the average acceptance length for each iteration. All these methods can be perceived as exact or approximate versions of sampling tokens from the conditional distribution $p(x_{t}|x_{<t})$. Therefore, their output perplexity is bounded by greedy decoding.
An orthogonal direction to improve speculative decoding is to improve the effectiveness of the small draft model. It is obvious that if more draft tokens are accepted, the overall inference speed will increase. BiLD (Kim et al., 2023) uses a model prediction alignment technique to better train the small model. Liu et al. (Liu et al., 2023) propose online speculative decoding to continually update the draft model based on observed input data. Instead, Rest (He et al., 2023) uses a retrieval model to produce draft tokens. An alternative way is to train additional heads in the large model to predict future tokens. Representative works include EAGLE (Li et al., 2024) and MEDUSA (Cai et al., 2024). Importantly, these works are orthogonal to speculative decoding techniques, including our proposed method. This orthogonality means that the improvements offered by more accurate draft tokens could be combined with our method for better effectiveness.
@@ 7 Conclusion
We introduce multi-token assisted decoding that significantly enhances output quality along with better time and energy efficiency. A distinctive aspect of our work is the exploration of speculative decoding’s impact on inference energy consumption, an often neglected area in existing studies. This research contributes not only a novel decoding approach but also valuable insights for optimizing LLM deployment in real-world applications where considerations of both quality and efficiency are crucial.

### v3

@@ 4 Energy Efficiency Analysis
Previous studies (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023; Sun et al., 2023) only focus on the speed of speculative decoding. However, an equally important consideration is energy consumption. To our knowledge, there is no existing work evaluating the impact of speculative decoding on inference energy consumption. Although MTAD and speculative decoding raise the number of FLOPs due to the involvement of a small auxiliary model and the rollback operation, they concurrently reduce the inference time and memory operations, which are key factors of GPU (or TPU) energy consumption (Allen & Ge, 2016; Chen et al., 2011). Consequently, it poses an open question regarding whether speculative decoding increases or decreases overall energy consumption.
@@
where $PW_{flop},PW_{mem}$ denote the power (energy/second) of a unit FLOP and memory operation, respectively, and $T_{flop},T_{mem}$ are the total time spent on these operations. When input batch size increases, $PW_{flop}$ rises until it reaches the power of maximum FLOPs, designated as $PW^{*}_{flop}$. $PW_{mem}$ is irrelevant to the input batch size because it only depends on the memory hardware.
@@
To determine the relative magnitude relationship between $PW_{flop}$ and $PW_{mem}$, we first point out the fact that GPU memory operations in LLM inference are dominated by accessing off-chip global memory. This consumes about 100×100$\times$100 × of energy compared to accessing on-chip shared memory (Jouppi et al., 2021). Because each multiprocessor on a GPU usually has 64KB of on-chip memory shared by multiple threads, but to store a single layer of LLM, say T5-11b (Raffel et al., 2020), requires about 1GB of memory. Moreover, Allen and Ge showed that doing a sequential read from off-chip memory consumes $20$-$30\%$ more power than running maximum FLOPs (Allen & Ge, 2016). So we have $PW_{mem}>PW^{*}_{flop}\geq PW_{flop}$. Notice that $PW^{*}_{flop}=PW_{flop}$ only if the batch size reaches the maximum parallelization capacity of GPUs. During multinomial sampling and speculative decoding, the batch size is generally small (Leviathan et al., 2023). So most of the computing power is not utilized (Leviathan et al., 2023), which means $PW_{mem}\gg PW_{flop}$.
In addition, previous studies have revealed that during LLM inference $T_{mem}\gg T_{flop}$ (Leviathan et al., 2023). Therefore, the energy induced by memory operations, i.e., $PW_{mem}T_{mem}$ dominates $E_{total}$. Since speculative decoding lowers $T_{mem}$ by reducing the number of runs of the large model, it should cut the inference energy consumption to a similar extent as it reduces time consumption.
To validate our hypothesis, we conducted an experiment to evaluate how batch size influences energy consumption during inference. We ran OPT-13b models on a Nvidia L40 GPUs with 48GB memory. Fixing the total number of runs of the large model while varying the input batch size $b\in\{1,2,4,8,16\}$ for each run, we measured time and energy cost. The details of energy measurement are illustrated in the Appendix D. Table 2 shows the results. As batch size doubles, although the number of FLOPs doubles, the energy consumption per run goes up slightly. This observation demonstrates that $PW_{mem}T_{mem}$ dominates $E_{total}$. Moreover, we measured the speed and energy consumption of running multinomial sampling with the large model and speculative decoding with OPT (125M, 13B) and Llama-2 (68M, 13B) models. The results, seen in Table 3, indicate that speculative decoding lowers the energy consumption and the time cost. This observation corroborates our claim to the energy efficiency of speculative decoding.
@@ 5 Experiments
Datasets and Models. In the main paper, we report results with three public datasets for evaluation: (1) Spider (Yu et al., 2018), MTBench (Zheng et al., 2023), and HumanEval (Chen et al., 2021). We use Llama-3-8B and Llama-3-8B-Instruct (Dubey et al., 2024) as target models, and Llama-3-1B and Llama-3-1B-Instruct as their draft models, respectively. We provide additional experiments with other datasets and model families in Appendix C.
Baselines. We compare our method with six speculative decoding methods, including four lossless decoding methods: vanilla speculative decoding (SpD(Lee et al., 2018; Chen et al., 2023), Spectr (Sun et al., 2023), SpecInfer (Miao et al., 2023), MCSS (Yang et al., 2024), and two lossy speculative decoding methods: BiLD (Kim et al., 2023) and typical decoding (Cai et al., 2024). All the baselines and our method utilize the same pair of draft and target models without any fine-tuning. For each method, we let it generate at most 128 tokens for each input and run it for $1,000$ seconds. All the methods are stochastic with top-$k$ and top-$p$ sampling with the temperature = 1. The details of the hyper-parameters (e.g., $k$ and $p$) and machine configurations of the experiments can listed in the Appendix D, E, and F.
@@
5.1 Performance of Multi-Token Joint Decoding
While most speculative decoding approaches focus on inference speed up, we want to design approaches that can also improve inference quality. We propose multi-token joint decoding (MTJD, Section 3.1) to accomplish the goal, due to its capability to achieve a lower perplexity and higher likelihood than single-token multinomial sampling. To validate that MTJD indeed betters output quality, we test MTJD (k=4) and standard multinomial sampling on Spider, MTBench, and HumanEval using the Llama-3 series models. We follow the same way introduced in Section 3.1 to implement MTJD. For this process, the higher the scores, the better the downstream performance. Under all settings, MTJD realizes the highest scores and lower perplexity. These results show a clear advantage for MTJD in terms of output quality.
Table 4: Performance comparison of single-token sampling and multi-token joint sampling. We use Llama-3.1-8B and Llama-3.1-8B-Instruct as target models, and Llama-3.2-1B and Llama-3.2-1B-Instruct as the draft models.
Llama-3 (8B,1B) Llama-3-Instruct (8B,1B)
  Spider  MTBench HumanEval   Spider  MTBench HumanEval
Single-token multinomial sampling
Score 22.0 3.40 15.9 36.0 4.11 28.0
PPL 2.58 2.40 2.09 2.23 1.91 1.85
Multi-token joint sampling
Score 52.5 3.77 36.6 60.5 4.40 49.4
PPL 1.16 1.32 1.26 1.18 1.27 1.15
5.2 Performance of Multi-Token Assisted Decoding
Next, we evaluate the efficiency and effectiveness of MTAD, an approximate algorithm that accelerates MTJD while preserving its downstream performance advantages. Table 5 presents the decoding speed, energy consumption, and downstream performance of various decoding algorithms across different datasets.
Efficiency Analysis.
We first observe that MTAD is the most efficient among all baselines in terms of both energy and time. On average, MTAD is 17.2% faster than the most efficient lossless baseline, MCSS, while consuming 6.1% less energy. Compared to lossy decoding algorithms, it attains 27.2% (15.1%) higher speed and 17.9% (14.1%) lower energy consumption than BiLD (typical decoding), respectively. Interestingly, despite utilizing only a single draft sequence, MTAD outperforms baselines that employ multiple draft sequences, such as Spectr, MCSS, and SpecInfer. In these methods, verification terminates immediately upon rejecting a token. In contrast, MTAD continues verification even after a rejection, searching for future tokens that may still pass. This mechanism results in a greater acceptance length per iteration than the baselines.
Downstream Performance Comparison.
Next, we compare the downstream performance of different decoding algorithms. Notice that while lossless decoding algorithms theoretically sample from the target distribution, they exhibit slight variations in downstream performance. This discrepancy arises because, despite preserving the original distribution, differences in the sampling process prevent them from generating identical sequences even when the random seed is fixed. Furthermore, we observe that lossy decoding algorithms can reach higher downstream performance at the expense of efficiency. This suggests that all lossy decoding methods can trade off efficiency for performance by adjusting verification strictness. Most notably, MTAD consistently achieves the highest downstream performance. On average, it surpasses lossless decoding algorithms by 25.0%, BiLD by 11.8%, and typical decoding by 9.9%. These results confirm our claim that MTAD offers superior effectiveness compared to conventional decoding methods that rely solely on single-token distributions.
Table 5: Comparison of different speculative decoding methods across various models and metrics. Bold indicates best values, underline indicates second-best.
Lossy Decoding Lossless Decoding Ours
BiLD Typical SpD Spectr SpecInfer MCSS MTAD
HumanEval
Llama-3-Instruct
tokens/s $\uparrow$ 17.4 21.7 22.2 23.8 22.8 23.7 24.8
J/token $\downarrow$ 10.0 8.1 7.8 7.8 7.9 7.8 7.6
pass@1 $\uparrow$ 37.8 35.9 32.9 32.9 31.0 32.0 38.4
Llama-3
tokens/s $\uparrow$ 19.6 22.5 22.2 24.4 22.5 23.8 25.6
J/token $\downarrow$ 9.7 8.9 8.9 8.9 8.1 7.9 7.6
pass@1 $\uparrow$ 19.5 20.0 15.9 16.0 17.7 17.0 22.0
Spider
Llama-3-Instruct
tokens/s $\uparrow$ 20.1 22.3 19.6 22.4 21.1 21.7 23.5
J/token $\downarrow$ 10.2 9.5 10.5 9.6 10.2 10.0 9.2
Acc $\uparrow$ 35.0 42.0 36.0 35.5 37.0 35.0 44.0
Llama-3
tokens/s $\uparrow$ 23.3 32.3 31.1 32.1 32.6 32.7 $\downarrow$ 33.3
J/token 8.2 7.9 7.5 7.1 8.1 8.0 7.8
Acc $\uparrow$ 30.5 29.5 21.5 23.0 21.5 24.0 35.0
MT-Bench
Llama-3-Instruct
tokens/s $\uparrow$ 25.9 23.4 26.0 26.2 26.3 26.8 $\downarrow$ 29.8
J/token $\downarrow$ 10.8 12.2 10.0 9.9 10.0 9.9 9.2
score $\uparrow$ 4.15 4.26 4.10 4.11 4.01 4.02 4.40
Llama-3
tokens/s $\uparrow$ 24.5 22.3 24.1 24.5 24.5 25.7 28.2
J/token $\downarrow$ 11.5 12.4 11.0 11.6 11.7 11.1 10.0
score $\uparrow$ 3.41 3.24 3.39 3.41 3.35 3.36 3.75
@@ 6 Related Work
Non-autoregressive decoding more closely resembles our work. It is first proposed by (Gu et al., 2017) to generate multiple tokens in parallel. That is, the model simultaneously predicts $p(x_{t+k}|x_{1:t})$ ($k=1,2,\ldots$). Subsequent studies further improved the performance of parallel decoding by incorporating additional information (Wang et al., 2019; Sun et al., 2019; Li et al., 2019) or employing additional iterations to refine predictions (Ghazvininejad et al., 2019; Lee et al., 2018; Guo et al., 2020). However, these works require continuous training of the model and generally either compromise the model effectiveness or require task-dependent techniques to attain a comparable performance (Kim et al., 2023).
Speculative Decoding. Speculative decoding was recently proposed in (Leviathan et al., 2023; Chen et al., 2023) as a way to accelerate LLM inference. Spectr (Sun et al., 2023) enhances speculative decoding by letting the small model generate multiple i.i.d. draft sequences. While speculative decoding and Spectr use the large model to verify all the tokens drafted by the small model, BiLD (Kim et al., 2023) only calls the large model when the probability output by the small model is below a pre-defined threshold $\tau_{1}$. The large model rejects a token if its negative log-likelihood is larger than threshold $\tau_{2}$. SpecInfer (Miao et al., 2023) utilizes one or multiple small models to generate a draft token tree to increase the average acceptance length for each iteration. MCSS (Yang et al., 2024) further strengthens SpecInfer via sampling without replacement. All these methods can be perceived as exact or approximate versions of sampling tokens from the conditional distribution $p(x_{t}|x_{<t})$. Therefore, their output perplexity is bounded by greedy decoding.
An orthogonal direction to boost speculative decoding is to improve the effectiveness of the small draft model. It is obvious that if more draft tokens are accepted, the overall inference speed will increase. BiLD (Kim et al., 2023) employs a model prediction alignment technique to better train the small model. Liu et al. (Liu et al., 2023) propose online speculative decoding to continually update the draft model based on observed input data. Instead, Rest (He et al., 2023) uses a retrieval model to produce draft tokens. An alternative way is to train additional heads in the large model to predict future tokens. Representative works include EAGLE (Li et al., 2024) and MEDUSA (Cai et al., 2024). Importantly, these works are orthogonal to speculative decoding techniques, including our proposed method. This orthogonality means that the improvements offered by more accurate draft tokens could be combined with our method for better effectiveness.
@@ 7 Conclusion
We introduce multi-token assisted decoding, a process that enhances output quality while improving time and energy efficiency A distinctive aspect of our work is the exploration of speculative decoding’s impact on inference energy consumption, an often neglected area in existing studies. This research contributes not only a novel decoding approach but also valuable insights for optimizing LLM deployment in real-world applications where considerations of both quality and efficiency are crucial.
8 Acknowlegement
This work was partially supported by NSF grants 2211557, 1937599, 2119643, 2303037, NSF 2312501, SRC JUMP 2.0 PRISM Center, NASA, Okawa Foundation, Amazon Research, Snapchat, and the CDSC industrial partners (https://cdsc.ucla.edu/partners/). The authors would also like to thank Marci Baun for editing the paper.